# Multi-scale Topology Optimization using Neural Networks

### v1

@@ 5 LIMITATIONS AND FUTURE WORK
In this study, our primary emphasis lies in showcasing our capability to generate interconnected microstructures amidst demanding design constraints. In the multi-scale topology optimization example, we rely on a mapping between the principal strain and the stiffness tensor. The use of isotropic material in coarse-scale optimization is also not a realistic representation of the stiffness of fine-scale structures. A wide variety of future work can be performed to expand the application of our proposed approach. Regarding the fine-scale optimization, we could change the objective function such that it becomes a stiffness tensor-constrained volume fraction minimization problem. This allows us to design the desired stiffness tensor from the coarse scale while optimizing fine-scale microstructure to achieve it.
The convergence of the multi-scale structure can be further explored. The examples in the paper, similar to full batch in machine learning, we simultaneously optimize all cells every epoch. However, a mini-batch can also be used such that a selection of cells is optimized. Future works could focus on exploring how to select the correct batch given the convergence of the problem to more effectively optimize multi-scale structures. Once the efficiency is improved, larger size problems with more cell count and extension to 3D can be explored. In a prior work [9], the authors explored using a conditioning field to speed up convergence on single-scale topology optimization problems. The use of a conditioning field can also be applied to multi-scale structures. One possible conditioning field is the rank-N laminates. Where the neural network can help with connecting the Rank-N laminates. In another prior work, [30], Physics Informed Neural Network (PINN) is used to directly optimize topology. It is worth exploring if PINN can also be used in multi-scale optimization, whether in coarse-scale designing stiffness tensors or in fine-scale designing microstructures.
Manufacturing constraints can also be explored. In prior works [31], the neural network-based topology optimization allows the computation of density gradients therefore overhang control can be applied for additive manufacturing. Overhang control can also be applied to microstructures such that each individual cell can be additively manufactured without suffering from local print failures. Multi-material, and multi-physics can also be integrated into the proposed approach.

### v2

@@ 5 LIMITATIONS AND FUTURE WORK
In this study, our primary emphasis lies in showcasing our capability to generate interconnected microstructures amidst demanding design constraints. The convergence of the multi-scale structure can be further explored. In a prior work [9], the authors explored using a conditioning field to speed up convergence on single-scale topology optimization problems. A conditioning field can also be used to apply to multi-scale structures. One possible conditioning field is the rank-N laminates. The neural network can help connect the Rank-N laminates. In another prior work, [35], Physics Informed Neural Network (PINN) is used to directly optimize topology. It is worth exploring if PINN can also be used in multi-scale optimization, whether in macro-scale designing stiffness tensors or in micro-scale designing microstructures.
Manufacturing constraints can also be explored. In prior works [36], the neural network-based topology optimization allows the computation of density gradients; therefore, overhang control can be applied for additive manufacturing. Overhang control can also be applied to microstructures such that each individual cell can be additively manufactured without suffering from local print failures. Multi-material and multi-physics can also be integrated into the proposed approach.
@@ Acknowledgment
We want to thank Professor Liwei Wang for insightful discussions and assistance in multi-scale topology optimization.